We are doing algebra equations I need help doing (n divided by 5 = 10)

Answer by jim_thompson5910(35256)

You need to isolate n. To do that, you undo the division by 5. Multiplication is the opposite of division

So you need to multiply both sides by 5


n/5 = 10

5*(n/5) = 5*10 ... Multiply both sides by 5.

5n/5 = 50

1n/1 = 50

1n = 50

n = 50

So the final answer is n = 50